OpenStudy (anonymous):

Solve. -3(7-6)-x=26-2(11 x+4)

OpenStudy (anonymous):

-21 +18 -x = 26 - 22x - 8 -3 -x = 14 -22x -3 -14 = -22x + x -17 = -21x x = -17/-21 = 17/21 better problem :)